# Cataclysmic Variables in Triples: Formation Models and New Discoveries

### v1

@@ 3.3 Binary Magnetic Braking
For close WD+MS binaries, both before and after the onset of mass transfer, magnetic braking (MB) is one of the main angular momentum loss mechanisms that drive their evolution (e.g., Verbunt & Zwaan, 1981; Rappaport et al., 1983). Since CV formation and evolution are sensitive to MB, we modify COSMIC to include the MB prescription from Rappaport et al. (1983). As we discuss with detail in Appendix C, the updated MB model leads to consistent evolution between the COSMIC binaries and MESA binary evolution for CVs, up to the period gap. Before changing the MB, many of the inner binaries in triples would merge almost instantly after mass transfer. Now, with the new MB prescription, binaries remain stably mass transferring for longer periods which matches MESA (Figure 11), as is consistent with observed populations of CVs (e.g., Rappaport et al., 1983). Note COSMIC merges the binary earlier than MESA predicts. Nevertheless, since our study focuses mainly on the properties at CV formation (i.e., first mass transfer), the subsequent CV evolution via COSMIC is not essential to our conclusions.
@@ 4.1 Outcomes of Population Synthesis
If we expand our search to include all types of non-compact object companions that are stably mass-transferring systems with a WD, such as giant and helium star donors, then we find systems. This means that there are times more mass transfer interactions between WD+evolved companion inner binaries than there are for WD+MS inner binaries444These triple-induced WD mass transfer rates are similar to those predicted for binaries in the Galactic Center (e.g., Stephan et al., 2019).. While these binaries would not be classified as CVs, they are expected to evolve into related types of systems, primarily symbiotic (red giant + WD) and AM CVn (helium star + WD) binaries. This large fraction of evolved donors transferring mass with a WD is not surprising, given that evolved companions have larger stellar radii, which increases the chance of interaction and the efficiency of tides to shrink the inner orbit. As we discuss further in Section 4.2, we expect the triple channel to preferentially produce CVs with more evolved donors and longer orbital periods than the standard binary channel for some fraction of the CVs. In Appendix D and Figure 12, we outline the characteristics of accreting WD systems with evolved donors, though the evolution of these systems in triples warrants further detailed study.
@@ 4.2 Formation of CVs in Triples
The last column of Figure 6 shows the triple configuration at the end of their evolution. We define the ‘end’ of a triples evolution as Gyr for the detached triples. If the inner binary became a CV, we instead plot the orbital parameters at the time of CV formation. If the inner binary merged and never was a CV, we define the final time as the time right before the merger event took place. Most orange and green points in this final panel are mergers, while the primary was an MS or RG. All of these have circular orbits just before the merger due to COSMIC’s instantaneous circularization upon mass transfer, although many originated from a highly eccentric inner orbit. Such merged systems are studied in detail in Shariat et al. (2024b). This panel also contains systems where a WD is accreting from a companion that is not a MS star; we highlight these systems in Figure 12 and Appendix D.
@@ 4.2.1 Channels Without a Common Envelope Phase
While the number of these CV outcomes here is relatively small, the fraction is consistent with rates of similar mass transfer channels reported in previous triple population studies (e.g., Toonen et al., 2016; Stephan et al., 2019; Shariat et al., 2023). Thus, simply applying Poisson error bars to these fractions would likely overestimate the uncertainty, and these results are robust. We chose not to simulate a larger population because we focused on modeling the full evolution for all 2000 triples to get an accurate and complete population, which required over 50,000 CPU hours. Approximately of these triples required more than 24 hours to converge, and some simulations required up to a week to complete555We note that triple integration techniques often impose a shorter integration time limit. Such a practice misses the majority of these systems..
@@ 5 Discussion
In this study, we do not self-consistently track eccentric mass transfer, which is especially important in three-body evolutionary scenarios. In our COSMIC binary simulations, the inner binary is circularized as soon as mass transfer begins, leading to qualitatively different outcomes compared to other eccentric mass transfer models (e.g., Sepinsky et al., 2009; Akira Rocha et al., 2024). Knigge et al. (2022), who identified a wide proper motion companion to the recurrent nova T Pyx, showed that highly eccentric mass transfer might play a role in such short-period recurrent novae. Future models with self-consistent eccentric mass transfer would be more accurate in following CV evolution, especially in triples where high eccentricities are common. In this work, we aim to broadly model the formation of CVs in triple, such that it may be applied on a population level. Note also that some of our outlined formation channels (Section 4.2) are similar to the proposed channels for T Pyx.
@@
White dwarf natal kicks from asymmetric mass loss on the asymptotic giant branch (AGB) would naturally widen the tertiary separations in the CV population. We do not consider the effects of mass loss kicks on the inner and outer orbits in our triple in our study. Imprints of WD kicks have been proposed in wide double WD binaries (El-Badry & Rix, 2018) and WD triples (Shariat et al., 2023). Additionally, for the CVs that form through CE, the mass loss experienced during the spiral-in phase is rapid compared to the outer binary period. In this case, the mass loss could provide an impulsive kick to the system, potentially widening or unbinding the tertiary star (e.g., Alcock et al., 1986; Toonen et al., 2017; Michaely & Perets, 2019; Igoshev et al., 2020). While the physics and timescales of CE evolution remain uncertain (e.g., Ivanova et al., 2013), previous studies report that the timescale for the CE phase is generally below yrs for WD+MS binaries (e.g., Ivanova et al., 2013; Igoshev et al., 2020). Therefore, rapid asymmetric mass loss can potentially generate a small impulsive kick within the triple during WD formation and the CE phase. This kick, although modest, can be comparable to the orbital velocity of CV tertiary companions since most are wide. This would unbind a fraction of the CV tertiaries, rendering them no longer bound to the CV and, therefore, not observable as CV triple. Either way, we expect many CVs with wide WD tertiaries to exist, though the faint WD companions render most of these currently unobservable. In Figure 8, we conservatively assume adiabatic mass loss, which widens the triple orbits in proportion to the mass lost from the system. In CV triples with WD tertiaries, there are two total WDs in the triple system, the tertiary was closer to the CV in the past, considering that mass loss has since widened the orbit. This would imply a stronger gravitational influence on the CV progenitor binary during the history of the triples evolution.
In Appendix E, we consider the impact of impulsive mass loss on resulting separations of CV-tertiaries. In summary, we find that such kicks shift the entire distribution by widening most binaries and ionizing several (Figure 13). However, we are hesitant to interpret these results given the small sample statistics, especially among the population that receives kicks. Nevertheless, assuming the presence of a mass-loss kick would suggest that even some isolated CVs formed through triple dynamics pathways, and their tertiary became unbound during the CV mass transfer evolution or white dwarf formation.
@@
Interestingly, the presence of a significant number of CVs with bound tertiary companions at separations exceeding au supports the idea that common envelope (CE) evolution does not always impart strong kicks, assuming that some fraction of these did experience a CE phase. At such large separations, the orbital velocity of the tertiary is km s-1, suggesting that any kick imparted due to rapid mass loss during the CE phase must be relatively small, likely no more than a few km s-1. This result can constrain models of CE evolution in support of scenarios where CE kicks are not large, such that wide triples can remain bound (e.g., Sandquist et al., 1998; Shiber et al., 2019; Chamandy et al., 2019). It may also be the case that all of these formed without a CE, though our simulations support that a fraction of them did.
@@
We note that the separation distribution of the simulations will be susceptible to the initial separation distribution (Figure 5). At the same time, the mass transfer outcomes are not highly sensitive to initial separations distributions (see, for example, Shariat et al., 2023). Still, if we only compare the CV triples to their wide binary counterparts, both of which are susceptible to similar systematics, we observe a robust excess of wide CV triples.
@@
The triple formation pathways proposed for CVs are remarkably similar to those proposed for the formation of black hole low-mass X-ray binaries (BH-LMXBs) (Naoz et al., 2016; Shariat et al., 2024a). Both classes of binaries involve a main-sequence (MS) donor star transferring mass onto a compact object— a white dwarf (WD) in the case of CVs and a black hole (BH) in the case of BH-LMXBs. In both cases, tertiary companions excite extreme eccentricities () to wide compact object+MS inner binaries, which can initiate mass transfer events before or after tidal circularization. Similar to the CVs, the BH-LMXB progenitors may be placed on these near-radial orbits either before the BH forms, in which case a CE phase may ensue, or after, in which case the mass transfer between the BH+donor star produces a BH-LMXB (Naoz et al., 2016; Shariat et al., 2024a). In contrast to a CV common envelope, the larger mass ratios between a BH progenitor and a low-mass MS companion might make successful envelope ejection more difficult in a BH-LMXB progenitor binary. Interestingly, a BH-LMXB in a hierarchical triple was recently observed (Burdge et al., 2024), confirming that BHs can form in triple environments. Overall, in both CVs and BH-LMXBs, triples provide a unique formation route in which the such can form without a CE phase. CVs are also more numerous than BH-LMXBs, allowing us to study their population demographics in more detail and distinguish between their various formation models. Insights from the CV population can potentially be used to learn more about BH-LMXB formation and evolution.
@@ 7 Acknowledgments
C.S. thanks, Ed Nathan, Natsuko Yamaguchi, and Samuel Whitebook for useful discussions. C.S. is supported by the Joshua and Beth Friedman Foundation Fund. S.N. acknowledges the partial support from NASA ATP 80NSSC20K0505 and from the NSF-AST 2206428 grant, as well as thanks Howard and Astrid Preston for their generous support.
@@ B.1 IGR J19308+0530
The Swift spectrum is best-fit by a 0.2 keV blackbody with erg cm-2 s-1, which corresponds to a 2-10 keV flux of erg cm-2 s-1 (Rodriguez et al., 2008). Ratti et al. (2013) identify an optical counterpart of spectral type F4V and find its radial velocity (RV) to vary with a period of days ( hrs); they associate this period with the orbital period of the binary. They measure a projected rotation velocity km s-1 and RV semi-amplitude km s-1. From this, they infer a mass ratio (see Equation B1; Horne et al., 1986). For typical F4 star with mass of M⊙, this would imply a faint companion mass of M⊙, which is consistent with a WD.
From Gaia, we learn that the system is pc away, and contains a third component au away. Given this distance, we can constrain the 2-10 keV X-ray luminosity to be erg s-1. Gaia also reports an RV for the tertiary of km s-1.
@@ Appendix E Impulsive Mass Loss
Throughout our study, we assume that mass loss in the inner binary leads to an adiabatic widening of the outer binary. While this has been generally accepted for slow, isotropic mass loss, the physical picture may change when the mass loss is sudden and anisotropic. The two major points in CV formation where this may occur are (1) when the AGB primary becomes a WD or (2) when the binary undergoes common envelope evolution, if it does. For the first case, previous studies suggest that WD binary and triple population show imprints of WD formation kicks (e.g., Fregeau et al., 2009; El-Badry & Rix, 2018; Shariat et al., 2023, 2024b). For rapid CE mass loss, the spiral-in phase may eject mass on timescales shorter than the tertiary star’s orbital period, potentially causing an impulsive kick (e.g., Toonen et al., 2017; Michaely & Perets, 2019; Igoshev et al., 2020).

### v2

@@ 3.3 Binary Magnetic Braking
For close WD+MS binaries, both before and after the onset of mass transfer, magnetic braking (MB) is one of the main angular momentum loss mechanisms that drive their evolution (e.g., Verbunt & Zwaan, 1981; Rappaport et al., 1983). Since CV formation and evolution are sensitive to MB, we modify COSMIC to include the MB prescription from Rappaport et al. (1983). We note that the Rappaport et al. (1983) magnetic braking prescription has several known limitations when compared to observations of both single and binary stars. Alternative models, such as saturated, disrupted, or boosted magnetic braking prescriptions (e.g., Ivanova & Taam, 2003; Matt et al., 2015; El-Badry et al., 2022; Belloni et al., 2024), may provide better agreement with the observed CV population. However, we do not explore these alternatives here, as our focus is primarily on the role of triple evolution prior to the common envelope phase. A comprehensive evaluation of different magnetic braking models and their impact on CV formation is beyond the scope of this work but is well warranted in future studies.
As we discuss with detail in Appendix C, the updated MB model leads to consistent evolution between the COSMIC binaries and MESA binary evolution for CVs, up to the period gap. Before changing the MB, many of the inner binaries in triples would merge almost instantly after mass transfer. Now, with the new MB prescription, binaries remain stably mass transferring for longer periods which matches MESA (Figure 11), as is consistent with observed populations of CVs (e.g., Rappaport et al., 1983). Note COSMIC merges the binary earlier than MESA predicts. Nevertheless, since our study focuses mainly on the properties at CV formation (i.e., first mass transfer), the subsequent CV evolution via COSMIC is not essential to our conclusions.
@@ 4.1 Outcomes of Population Synthesis
If we expand our search to include all types of non-compact object companions that are stably mass-transferring systems with a WD, such as giant and helium star donors, then we find systems. This means that there are times more mass transfer interactions between WD+evolved companion inner binaries than there are for WD+MS inner binaries666These triple-induced WD mass transfer rates are similar to those predicted for binaries in the Galactic Center (e.g., Stephan et al., 2019).. While these binaries would not be classified as CVs, they are expected to evolve into related types of systems, primarily symbiotic (red giant + WD) and AM CVn (helium star + WD) binaries. This large fraction of evolved donors transferring mass with a WD is not surprising, given that evolved companions have larger stellar radii, which increases the chance of interaction and the efficiency of tides to shrink the inner orbit. As we discuss further in Section 4.2, we expect the triple channel to preferentially produce CVs with more evolved donors and longer orbital periods than the standard binary channel for some fraction of the CVs. In Appendix D and Figure 12, we outline the characteristics of accreting WD systems with evolved donors, though the evolution of these systems in triples warrants further detailed study.
@@ 4.2 Formation of CVs in Triples
The last column of Figure 6 shows the triple configuration at the end of their evolution. We define the ‘end’ of a triples evolution as Gyr for the detached triples. If the inner binary became a CV, we instead plot the orbital parameters at the time of CV formation. If the inner binary merged and never was a CV, we define the final time as the time right before the merger event took place. Most orange and green points in this final panel are mergers, while the primary was an MS or RG. All of these have circular orbits just before the merger due to COSMIC’s instantaneous circularization upon mass transfer, although many originated from a highly eccentric inner orbit. Such merged systems are studied in detail in Shariat et al. (2024). This panel also contains systems where a WD is accreting from a companion that is not a MS star; we highlight these systems in Figure 12 and Appendix D.
@@ 4.2.1 Channels Without a Common Envelope Phase
While the number of these CV outcomes here is relatively small, the fraction is consistent with rates of similar mass transfer channels reported in previous triple population studies (e.g., Toonen et al., 2016; Stephan et al., 2019; Shariat et al., 2023). Thus, simply applying Poisson error bars to these fractions would likely overestimate the uncertainty, and these results are robust. We chose not to simulate a larger population because we focused on modeling the full evolution for all 2000 triples to get an accurate and complete population, which required over 50,000 CPU hours. Approximately of these triples required more than 24 hours to converge, and some simulations required up to a week to complete777We note that triple integration techniques often impose a shorter integration time limit. Such a practice misses the majority of these systems..
@@ 5 Discussion
In this study, we do not self-consistently track eccentric mass transfer, which is especially important in three-body evolutionary scenarios. In our COSMIC binary simulations, the inner binary is circularized as soon as mass transfer begins, leading to qualitatively different outcomes compared to other eccentric mass transfer models (e.g., Sepinsky et al., 2009; Hamers & Dosopoulou, 2019; Glanz & Perets, 2021; Akira Rocha et al., 2024). Knigge et al. (2022), who identified a wide proper motion companion to the recurrent nova T Pyx, showed that highly eccentric mass transfer might play a role in such short-period recurrent novae. Note also that some of our outlined formation channels (Section 4.2) are similar to the proposed channels for T Pyx. Future models with self-consistent eccentric mass transfer would be more accurate in following CV evolution, especially in triples where high eccentricities are common. Future models could also explore different angular momentum loss and mass transfer stability prescriptions, which may change the fraction of systems in a population that evolve into CVs (e.g., Belloni et al., 2018). One relevant change for modeling CVs is incorporating consequential angular momentum loss (CAML) (King & Kolb, 1995; Schreiber et al., 2016). In this work, however, our aim is to broadly model CV formation pathways in triple systems up to the onset of the first mass transfer phase.
@@
White dwarf natal kicks from asymmetric mass loss on the asymptotic giant branch (AGB) would naturally widen the tertiary separations in the CV population. We do not consider the effects of mass loss kicks on the inner and outer orbits in our triple in our study. Imprints of WD kicks have been proposed in wide double WD binaries (El-Badry & Rix, 2018) and WD triples (Shariat et al., 2023). Additionally, for the CVs that form through CE, the mass loss experienced during the spiral-in phase is generally rapid compared to the orbital timescales. In this case, the mass loss could provide an impulsive kick to the system, potentially widening or unbinding the tertiary star (e.g., Alcock et al., 1986; Toonen et al., 2017; Michaely & Perets, 2019a; Igoshev et al., 2020). Early works (Michaely & Perets, 2019a) proposed using wide companions to constrain CE timescales and mass loss, which was later pursued statistically by Igoshev et al. (2020). While the physics and timescales of CE evolution remain uncertain (e.g., Ivanova et al., 2013), previous studies report that the timescale for the CE phase is generally below yrs for WD+MS binaries (e.g., Ivanova et al., 2013; Michaely & Perets, 2019a; Igoshev et al., 2020). Therefore, rapid or asymmetric mass loss can potentially generate a small impulsive kick within the triple during WD formation and/or the CE phase, if one occurs. This kick, although modest, is often comparable to the orbital velocity of CV tertiary companions since most tertiaries are wide. Such a kick would unbind a fraction of the CV tertiaries, rendering them no longer bound to the CV and thereby no longer observable as CV triple. Note that since the tertiary was bound up until WD formation (or the first CE phase), secular three-body dynamics could still be influential to the system’s pre-mass transfer evolution. With or without WD kicks, we expect CVs with wide WD tertiaries to exist, with most faint WD companions being difficult to detect with Gaia.
In Figure 8, we conservatively assume the effects of adiabatic mass loss, which widens the triple orbits in proportion to the mass lost from the system (see also Kratter & Perets, 2012; Shappee & Thompson, 2013; Michaely & Perets, 2014, for early considerations of mass-loss into secular evolution). There also exists a mass loss regime between the adiabatic and impulsive approximations, in which both the semi-major axis and eccentricity can change. We do not expect this scenario to change our results significantly since considering both limiting cases (i.e., adiabatic and impulsive) did not change our main conclusions (Appendix E). In CV triples with WD tertiaries, there are two total WDs in the triple system suggesting two episode of mass loss in the triple. Therefore, the tertiary was likely closer to the CV in the past considering that the orbits widened during mass loss, implying a stronger gravitational influence on the CV progenitor binary during the history of the triples evolution.
@@
In Appendix E, we consider the impact of impulsive mass loss on resulting separations of CV-tertiaries. The physical mechanism for this is kept agnostic, but may apply to WD natal kicks or CE kicks due to rapid and/or asymmetric mass loss. In summary, we find that such kicks shift the entire distribution by widening most binaries and ionizing some (Figure 13). Assuming mass-loss kicks would suggest that an even larger fraction of isolated CVs formed through triple dynamics pathways, and their tertiary became unbound during the CV mass transfer evolution or white dwarf formation. Interestingly, the presence of a significant number of CVs with bound tertiary companions at separations exceeding au supports the idea that common envelope (CE) evolution does not always impart strong kicks, since some fraction of these systems likely experienced a CE phase. At such large separations, the orbital velocity of the tertiary is km s-1, suggesting that any kick imparted due to rapid mass loss during the CE phase must be relatively small, likely no more than a few km s-1. This result supports models where CE ‘kicks’ are not large or where the CE phase is not too rapid (relative to orbital timescales), such that wide triples can remain bound post-CE (e.g., Sandquist et al., 1998; Shiber et al., 2019; Chamandy et al., 2019; Michaely & Perets, 2019a; Igoshev et al., 2020). It may also be the case that all of these formed without a CE, though our simulations support that a fraction of them did. Another scenario where CVs can form without a CE phase is in dense cluster environments, where strong dynamical encounters and exchange can occur (Knigge, 2012; Belloni et al., 2016, 2017; Belloni & Rivera Sandoval, 2021; Kremer et al., 2020). However, recent observations suggest that dynamical interactions in globular clusters cannot efficiently form CVs (Rivera Sandoval et al., 2024).
@@
We note that the separation distribution of the simulations will be susceptible to the initial separation distribution (Figure 5). At the same time, the mass transfer outcomes are not highly sensitive to initial separations distributions (see, for example, Shariat et al., 2023). Even if we only compare the CV triples to their wide binary counterparts, both of which are susceptible to similar systematics, we still observe a robust excess of wide CV triples.
@@
The triple formation pathways proposed for CVs are remarkably similar to those proposed for the formation of black hole low-mass X-ray binaries (BH-LMXBs) (Naoz et al., 2016; Shariat et al., 2025). Both classes of binaries involve a main-sequence (MS) donor star transferring mass onto a compact object— a white dwarf (WD) in the case of CVs and a black hole (BH) in the case of BH-LMXBs. In both cases, tertiary companions excite extreme eccentricities () to wide compact object+MS inner binaries, which can initiate mass transfer events before or after tidal circularization. Similar to the CVs, the BH-LMXB progenitors may be placed on these near-radial orbits either before the BH forms, in which case a CE phase may ensue, or after, in which case the mass transfer between the BH+donor star produces a BH-LMXB (Naoz et al., 2016; Shariat et al., 2025). In contrast to a CV common envelope, the larger mass ratios between a BH progenitor and a low-mass MS companion might make successful envelope ejection more difficult in a BH-LMXB progenitor binary. Observationally, testing the common envelope scenario may be possible by using molecular abundance ratios as tracers (e.g., Sarna et al., 1995; Dhillon et al., 2002). Interestingly, a BH-LMXB in a hierarchical triple was recently observed (Burdge et al., 2024), confirming that BHs can form in triple environments. Overall, in both CVs and BH-LMXBs, triples provide a unique formation route in which the such can form without a CE phase. CVs are also more numerous than BH-LMXBs, allowing us to study their population demographics in more detail and distinguish between their various formation models. Insights from the CV population can potentially be used to learn more about BH-LMXB formation and evolution.
@@ 7 Acknowledgments
We thank the referee for constructive comments on the manuscript. C.S. thanks Ed Nathan, Natsuko Yamaguchi, and Samuel Whitebook for useful discussions. C.S. is supported by the Joshua and Beth Friedman Foundation Fund. S.N. acknowledges the partial support from NASA ATP 80NSSC20K0505 and from the NSF-AST 2206428 grant, as well as thanks Howard and Astrid Preston for their generous support. This research was supported by NSF grant AST-2307232.
@@ B.1 IGR J19308+0530
The Swift spectrum is best-fit by a 0.2 keV blackbody with erg cm-2 s-1, which corresponds to a 2-10 keV flux of erg cm-2 s-1 (Rodriguez et al., 2008). Ratti et al. (2013) identify an optical counterpart of spectral type F4V and find its radial velocity (RV) to vary with a period of days ( hrs); they associate this period with the orbital period of the binary. They measure a projected rotation velocity km s-1 and RV semi-amplitude km s-1. From this, they infer a mass ratio (see Equation B1; Horne et al., 1986). For a typical F4 star with a mass of M⊙, this would imply a faint companion mass of M⊙, which is consistent with a WD.
From Gaia, we learn that the system is pc away and contains a third component au away. Given this distance, we can constrain the 2-10 keV X-ray luminosity to be erg s-1. Gaia also reports an RV for the tertiary of km s-1.
@@ Appendix E Impulsive Mass Loss
Throughout our study, we assume that mass loss in the inner binary leads to an adiabatic widening of the outer binary. While this has been generally accepted for slow, isotropic mass loss, the physical picture may change when the mass loss is sudden and anisotropic. The two major points in CV formation where this may occur are (1) when the AGB primary becomes a WD or (2) when the binary undergoes common envelope evolution, if it does. For the first case, previous studies suggest that WD binary and triple population show imprints of WD formation kicks (e.g., Fregeau et al., 2009; El-Badry & Rix, 2018; Shariat et al., 2023, 2024). For rapid CE mass loss, the spiral-in phase may eject mass on timescales shorter than the tertiary star’s orbital period, potentially causing an impulsive kick (e.g., Toonen et al., 2017; Michaely & Perets, 2019b; Igoshev et al., 2020).